Week of 061127

Open Actions from last week:

Chair: H.Renshall

Gmod: M.Barosso

Smod: J.van Eldik


Log: CMS Castor2 garbage collection failed over the weekend - OB had to trigger manual runs. Also a user provoking many copies from the CSA06 pool to the default pool caused disk server high loads leading to many broken disk mirrors. This is being followed up.

New Actions: Start multi-VO test to IN2P3. ALICE and CMS first, GSSDATLAS next week. New IN2P3 end points were sent to experiments.

Discussion: No outbound FTS traffic over weekend. ALICE are changing Alien version and should resume transfers soon.


Log: Castor2 GSSDLHCB transaction logs got corrupted - fixed by JvE - down 1 hour.

New Actions:

Discussion: Thousands of GSSDATLAS transfer failures to BNL due to incorrect directory permissions at BNL. CASTORCMS software upgrade scheduled for tomorrow 09.30-10.30 (should fix garbage collection problems).


Log: Network problems at CERN.

New Actions:

  • Castor c2cms upgrade today - 9.30 (Jan)
  • Check in opm grid templates for reference to prod-grid-cern-admins (James)
  • lcg-mon-gridftp - new version of rgma-python deployed on dteam cluster for testing. Q:(Jan) How was this done - didn't see a profile change ?



New Actions:

Discussion: SRM problems at IN2P3. Multi-VO test suspended.



New Actions:

Discussion: Check with IN2P3 Monday morning to see if can resume multi-VO tests.

Edit | Attach | Watch | Print version | History: r5 < r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r5 - 2007-02-02 - FlaviaDonno
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback